OpenStudy (anonymous):

Dale is making a small table in the shape of a right triangle. If two of the sides will measure about 12 inches each, how many inches should the longest side (c) measure? Use the appropriate formula and round your answer to the nearest tenth. 4.9 in 6.9 in 17.0 in 24.0 in

OpenStudy (anonymous):

ok since two of the sides are 12 inches each, that means the triangle is a 45-45-90 degree triangle, which allows us to use the pythagorean theorem. a^2+b^2=c^2 12^2+12^2=c^2 144+144=c^2 288=c^2 c=16.97 in. or in this case, rounded to 17 inches.